# Put CASH on Bandits: A Max K-Armed Problem for Automated Machine Learning

### v1

@@ Abstract
The Combined Algorithm Selection and Hyperparameter optimization (CASH) is a challenging resource allocation problem in the field of AutoML. We propose MaxUCB, a max $k$-armed bandit method to trade off exploring different model classes and conducting hyperparameter optimization. MaxUCB is specifically designed for the light-tailed and bounded reward distributions arising in this setting and, thus, provides an efficient alternative compared to classic max $k$-armed bandit methods assuming heavy-tailed reward distributions. We theoretically and empirically evaluate our method on four standard AutoML benchmarks, demonstrating superior performance over prior approaches.
@@ 1 Introduction
The performance of machine learning (ML) solutions is highly sensitive to the choice of algorithms and their hyperparameter configurations which can make finding an effective solution a challenging task. AutoML aims to reduce this complexity and make ML more accessible by automating these critical choices (Hutter et al., 2019; Baratchi et al., 2024). For example, Hyperparameter optimization (HPO) methods focus on finding well-performing hyperparameter settings given a resource constraint, such as an iteration count or a time limit. However, in practice, it is often unclear which ML model class would perform best on a given dataset. The problem of jointly searching the model class and the appropriate hyperparameters has been coined CASH, Combined Algorithm Selection and Hyperparameter optimization (Thornton et al., 2013). As a prime example, on tabular data, a ubiquitous data modality (van Breugel and van der Schaar, 2024), the state-of-the-art ML landscape covers classic ML methods, ensembles of gradient-boosted decision trees, and modern deep learning approaches (Kadra et al., 2021; Gorishniy et al., 2021, 2024; McElfresh et al., 2023; Kohli et al., 2024; Hollmann et al., 2023; Holzmüller et al., 2024).
@@
Figure 1: (Left) Running HPO for the best-performing model class (brown line) outperforms the combined search (black line), which runs HPO for the joint algorithm and hyperparameter configuration space. (Moddile) The empirical performance of model classes has an irregular distribution. It is left-skewed, and a higher mean does not necessarily correspond to a higher maximum. (Right) Illustration of our approach at time step $t$.
@@
A popular approach to address the CASH problem is to use categorical and conditional hyperparameters to run HPO directly on the combined hierarchical search space of models and hyperparameters. AutoML systems use this approach, which we call combined search, to search well-performing ML pipelines (Thornton et al., 2013; Feurer et al., 2015; Komer et al., 2014; Kotthoff et al., 2017; Feurer et al., 2022), but HPO remains inefficient in high-dimensional and hierarchical search spaces. A naive solution to address the scalability limitation is to run HPO independently for the smaller search spaces of each ML model class and then compare the found solutions. However, this solution often exceeds available computational resources and does not scale well with an increasing number of ML models. Figure 1 (left) illustrates the difference between searching each space individually (colored lines) and combined search (black line) on an exemplary dataset.
@@
To leverage the efficiency of HPO in low dimensions, we use a Multi-Armed Bandit (MAB) method (Lattimore and Szepesvári, 2020) to dynamically allocate our budget. As shown in Figure 1 (left), each time the bandit strategy pulls an arm, it runs one iteration of HPO to evaluate a new configuration, resulting in a loss. The negative of the loss is used as reward feedback for the bandit algorithm. This approach is known as decomposed CASH (Hoffman et al., 2014; Liu et al., 2019).
@@
While most classical MAB problems aim to maximize the average rewards over time, the goal of the bandit algorithm for decomposed CASH should be to maximize the maximum reward observed over time: as illustrated in Figure 1 (middle), tuning the LightGBM (red) will eventually outperform CatBoost. This goal aligns with Max $K$-Armed Bandit (MKB) problems (Carpentier and Valko, 2014; Achab et al., 2017; Baudry et al., 2022), often referred to as Extreme Bandits.
In the context of AutoML, the time horizon is limited, making efficiency crucial. Prior work on decomposed CASH found that state-of-the-art MKB algorithms are not sufficiently sample-efficient in practice (Hu et al., 2021; Balef et al., 2024). Additionally, Nishihara et al. (2016) argued that the parametric assumptions derived from extreme value theory are not applicable in the context of HPO and stated that “the question of what realistic assumptions are likely to hold in practice for hyperparameter optimization is an important question”.
We precisely address this open question through a thorough statistical analysis of the empirical reward distributions of HPO tasks. Our data-driven approach reveals that common assumptions used in MKB problems do not hold in this context, and we define new ones for MaxUCB, an MKB algorithm for bounded light-tailed reward distributions. Concretely, our contributions are as follows:
  1. 1.
    We study the empirical reward distributions in HPO tasks and observe that they follow negatively skewed bounded distributions (left-skewed or left-tailed reward distributions).
  2. 2.
    We propose an appropriate MKB algorithm, MaxUCB, for this newly defined setting.
  3. 3.
    We provide a regret bound and theoretical analysis for MaxUCB, showing that the number of sub-optimal trials grows sub-linearly with time.
  4. 4.
    Finally, we demonstrate superior empirical performance on four popular AutoML benchmark tasks where we compare MaxUCB with state-of-the-art MAB algorithms developed for AutoML, and combined search.
@@
Our main contribution is a state-of-the-art algorithm for decomposed CASH based on a novel extreme bandit algorithm we call MaxUCB (Algorithm 1). We demonstrate the performance of our method on four benchmarks, which highlight the relevance of our assumptions for a wide variety of CASH problems. We analyze the theoretical performance of MaxUCB (Theorem 4.2) through high-probability regret bounds that also justify our novel choice of exploration bonus for the type of distributions resulting from the CASH problem. Importantly, our objective is rather to propose a practical algorithm with good empirical performance on CASH rather than a general multi-purpose bandit algorithm, so our guarantees hold under carefully crafted assumptions that resolve previously open questions (Nishihara et al., 2016).
@@ 2 Solving CASH using Bandits
For our approach, we study the decomposed variant (Hoffman et al., 2014; Liu et al., 2019), which yields the following bilevel optimization problem depicted in Figure 1 (right): at the upper level, we aim to find the overall best-performing ML model $A^{(i^{*})}$ by selecting model $A^{(i)}\in\mathcal{A}$ iteratively, and at the lower level, we aim to find the best-performing configuration $\boldsymbol{\lambda^{*}}\in\boldsymbol{\Lambda}^{(i)}$ for the selected model $A^{(i)}$. Formally,
$A^{(i^{*})}\in\operatorname*{arg\,min}_{A^{(i)}\in\mathcal{A}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}^{*}},{\mathbb{D}}),\quad\text{s.t.}\quad\boldsymbol{\lambda}^{*}\in\operatorname*{arg\,min}_{\boldsymbol{\lambda}\in\boldsymbol{\Lambda^{(i)}}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}},{\mathbb{D}})$ (2)
The right-hand side of Equation 2 in the lower level can be efficiently addressed by existing iterative HPO methods such as Bayesian optimization (BO) (Jones et al., 1998; Garnett, 2022), which has been demonstrated to perform well in practical settings (Snoek et al., 2012; Chen et al., 2018; Cowen-Rivers et al., 2022). BO fits a surrogate model and uses an acquisition function to find a promising configuration to evaluate next. The left-hand side of Equation 2, corresponding to the upper level, has to trade off exploration and exploitation of HPO runs, a typical MAB problem as already noticed in prior work (Cicirello and Smith, 2005; Streeter and Smith, 2006a; Nishihara et al., 2016; Metelli et al., 2022).
We denote $\boldsymbol{\lambda}_{t}$ the configuration proposed by the HPO method in the lower level and $r_{i,t}$ to be the feedback to arm $i$ obtained by evaluating $\boldsymbol{\lambda}_{t}\in\boldsymbol{\Lambda}^{(i)}$ at time step $t\leq T$. To be consistent with the bandit literature, we consider the negative loss as the reward and maximize it:
@@
The goal is then to find the best-performing algorithm $A^{(i^{*})}$ and its optimal configuration $\boldsymbol{\lambda}^{*}$ under a budget constraint $T$. This can be framed as minimizing the regret $R(T)$ with $I_{t}$ being the selected arm at time $t$ by:
@@
Instead of using MKB algorithms to directly address Equation 4, closely related prior works focus on alternative methods. For example, Hu et al. (2021) proposed and analyzed the Extreme-Region Upper Confidence Bounds (ER-UCB) algorithm maximizing the extreme-region of the feedback distribution, assuming Gaussian rewards.
More recently, Balef et al. (2024) have shown that existing MKB algorithms underperform when applied to decomposed CASH, and they proposed methods for maximizing the quantile values instead of the maximum value.
As another alternative method, Li et al. (2020) framed the CASH problem as a Best Arm Identification (BAI) task and introduced the Rising Bandits algorithm (Li et al., 2020).
This MAB method assumes that the reward function for each arm increases with each pull, following a rested bandit model with non-decreasing payoffs (Heidari et al., 2016) (which has been shown to have linear regret when the reward increment per pull exceeds a threshold (Metelli et al., 2022)). Rising Bandits can be used for our setting using the maximum observed performance of the HPO history as the reward. However, this algorithm assumes deterministic rewards and increasing concave reward functions. To weaken this assumption, Li et al. (2020) introduced a hyperparameter to increase initial exploration. Mussi et al. (2024) further weakened this assumption by assuming that the moving average of the rewards is an increasing concave function.
Generally, in BAI approaches, the main goal is to identify the arm with the highest mean reward. The learner can freely use its resources to explore the arms and get a noisy observation of their mean value, but must eventually stop and output a single arm. This objective differs from ours and is ill-suited for CASH because the best algorithm has the best achievable performance rather than the best performance on average over hyperparameters.
@@ 3 Data Analysis of HPO Tasks
Here, we analyze and empirically examine our problem’s reward distributions. In our setup, the reward corresponds to the performance of the hyperparameter configurations obtained by running HPO. To illustrate the reward distributions, for each ML model class (corresponding to arms in our setup), we run $T=200$ iterations of HPO, with $32$ repetitions (each using a different seed) on a varying number of datasets on four AutoML benchmarks (see Appendix D.2 and  C.1).
We begin by analyzing our reward characteristics through the survival function $G$ of the distribution $d$ defined as:
$\forall x\in\mathbb{R},\quad G(x)=P_{X\sim d}(X\geq x).$
Figure 2 shows the average empirical survival function of all observed performances (normalized between 00 and $1$ for each task) for each arm. We rank model classes (i.e., arms) based on their best performance per dataset and report results on two benchmarks. We observe that the reward distributions are
@@
Figure 2: (Left two plots) The average empirical survival function of the reward distribution per arm ranked per dataset. Thin lines correspond to segments of the reward sequence and show the distribution change over time. (Right two plots) The average empirical survival function per dataset for the best and worst arm. Thin lines correspond to individual datasets.
@@
Suppose Assumption 3.2 holds. Then, there exists $L,U\geq 0$ such that the survival function $G$ can be bounded near $b$ by two linear functions.
Lemma 3.3 (proof in Appendix B.1) provides a way to characterize the shape of any bounded distribution near its maximum value through distribution-dependent constants, $L$ and $U$
(see Appendix C.2 for more details and a visualization). Intuitively, it quantifies the behavior of the distribution of the ML model performance in a given hyperparameter space near the optimum. A large value of $L$ indicates a steep drop in the survival function near the maximum, while a small value of $U$ leads to a more gradual decay of the survival function and conversely (see Figure 3).
@@
Figure 4: $L$ and $U$ values across individual datasets.
@@
Our empirical analysis shows that existing MKB algorithms’ assumptions about reward distributions are either overly broad or misaligned with the CASH problem. We identify two key differences between our assumptions and those made by current MKB algorithms. First, the sub-optimality gap makes the regret definition (Equation 4) for bounded distributions meaningful. Without the existence of a nonzero gap, the regret definition fails since any policy consistently selecting a single arm can achieve zero regret, as Nishihara et al. (2016) pointed out with Bernoulli distributions. Second, our reward distributions differ from those used in existing MKB algorithms. Lemma 3.3 characterizes the shape of the distribution, with a higher $L$ ensuring more mass near extreme values, making extreme values easier to estimate. In our case, values for $L$ are mostly higher than $1$ while for heavy-tailed distributions, commonly used as the basis in existing MKB algorithms, are close to 00 (see the rightmost plot in Figure 3). In general, our analysis shows that considering the right range of $L$ values unlocks the problem raised by Nishihara et al. (2016), who focused on constructing counterexamples through the distributions that have absurd values for $L$. Thus, Theorem $11$, “no policy can be guaranteed to perform asymptotically as well as an oracle that plays the single best arm for a given time horizon.”, which means any policy needs to explore all arms for budget T, does not apply here.
@@ 4 MaxUCB
Our exploration bonus $(\frac{\alpha\log(t)}{n})^{2}$ deviates from typical UCB literature due to faster concentration of maximum values in bounded distributions. The probability of bad events (see Equation 31 in Appendix B.3) can be written as:
$P(\text{Bad events})\leq e^{-n\sqrt{C(n)}}+nC(n).$ (6)
Then, setting $C(n)=1/n^{2}$ minimizes the probability of bad events: the first term becomes independent of $n$, while the second term decreases with $n$. Furthermore, MaxUCB uses $\alpha\geq 0$ to control the exploration-exploitation trade-off; a higher $\alpha$ leads to more exploration.
@@
1:$\alpha$(exploration parameter), $T$(time horizon), $K$(arms)
@@
3:    Pull arm $i$, set $n_{i}\leftarrow 1$, observe reward $r_{i,1}$triangleright Evaluating default configuration
@@
7:        Update policy $U_{i}=\max{(r_{i,1},...,r_{i,n_{i}})}+(\frac{\alpha\log(t)}{n_{i}})^{2}$
@@
Analysis of MaxUCB.
We first show a regret decomposition result specific to max-$K$ armed bandits (Proposition 4.1) that directly relates the regret definition in equation 4 with the number of suboptimal trials.
@@
(Upper Regret Bound) the upper regret bound up to time $T$ is related to the number of times sub-optimal arms are pulled,
@@
Where $N_{i}(T)=\mathbbm{E}(\sum\nolimits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, arm $i^{*}$ is the optimal arm and $b_{i}$ is the upper bound on the reward of arm $i$ as given by Assumption 3.2, i.e., the reward of arm $i$ lies within the interval $[a_{i},b_{i}]$.
@@
The result of Theorem 4.2 (proof provided in Appendix B.3) highlights that MaxUCB primarily leverages two key properties: the sub-optimality gap $\Delta_{i}$ and the shape of the distribution, as defined in Lemma 3.3. Specifically, the performance improves with a larger value for $\Delta_{i}$ (the sub-optimality gap) and higher values of $L_{i^{*}}$ ($L$ for the optimal arm), which means that samples drawn from the distribution of the optimal arm are likely to be close to the extreme values. Additionally, smaller values of $U_{i}$ ($U$ for a sub-optimal arm $i$), which means it is less likely to draw samples close to the extreme values, reduce the number of selecting a sub-optimal arm $i$, thus enhancing overall performance. For our task, as is shown in Figure 3, the values for $L_{i^{*}}$ are higher than $1$ and $U_{i}$ less than $10$ in most cases, yielding high empirical performance. We compare the number of pulls observed in our experiments with the expected values based on our theoretical analysis in Appendix E.2, showing that our analysis is not loose.
However, it is essential to note that, in general, finding the optimal arm is challenging if $\Delta_{i}$ is close to zero or $L_{i^{*}}$ is very small. We assess the performance of MaxUCB on synthetic tasks in Appendix E.4, showing that the performance of MaxUCB deteriorates if the characteristics deviate from our assumptions.
Finally, we provide a regret upper bound that combines the decomposition in Proposition 4.1 with the individual upper bounds above. Indeed, this requires finding a parameter $\alpha$ that resolves a trade-off between the arms and minimizes the total upper bound. This is done in Corollary 10, whose proof is immediate.
@@
If $L_{i^{*}}$, $\min\limits_{i\neq i^{*}}(\Delta_{i})$ and $T$ are known in advance, the exploration parameter $\alpha$ can be chosen as:
$\alpha=\frac{1}{4L_{i^{*}}\sqrt{\min\limits_{i\neq i^{*}}\Delta_{i}}}\bigg{(}1-\frac{2\log(\log(T))}{\log(T)}\bigg{)}.$ (9)
Then, the total regret is bounded by:
$R(T)\leq\mathcal{O}\bigg{(}\frac{K\log T}{\sqrt{T}}\max\limits_{i\leq K}b_{i}\bigg{)}.$ (10)

### v2

@@ Abstract
The Combined Algorithm Selection and Hyperparameter optimization (CASH) is a challenging resource allocation problem in the field of AutoML. We propose MaxUCB, a max $k$-armed bandit method to trade off exploring different model classes and conducting hyperparameter optimization. MaxUCB is specifically designed for the light-tailed and bounded reward distributions arising in this setting and, thus, provides an efficient alternative compared to classic max $k$-armed bandit methods assuming heavy-tailed reward distributions. We theoretically and empirically evaluate our method on four standard AutoML benchmarks demonstrating superior performance over prior approaches. We make our code and data available at https://github.com/amirbalef/CASH_with_Bandits.
@@ 1 Introduction
The performance of machine learning (ML) solutions is highly sensitive to the choice of algorithms and their hyperparameter configurations which can make finding an effective solution a challenging task. AutoML aims to reduce this complexity and make ML more accessible by automating these critical choices (Hutter et al., 2019; Baratchi et al., 2024). For example, Hyperparameter optimization (HPO) methods focus on finding well-performing hyperparameter settings given a resource constraint, such as an iteration count or a time limit. However, in practice, it is often unclear which ML model class would perform best on a given dataset (Bischl et al., 2025). The problem of jointly searching the model class and the appropriate hyperparameters has been coined CASH, Combined Algorithm Selection and Hyperparameter optimization (Thornton et al., 2013). As a prime example, on tabular data, a ubiquitous data modality (van Breugel and van der Schaar, 2024), the state-of-the-art ML landscape covers classic ML methods, ensembles of gradient-boosted decision trees and modern deep learning approaches (Kadra et al., 2021; Gorishniy et al., 2021, 2024; McElfresh et al., 2023; Kohli et al., 2024; Hollmann et al., 2023; Holzmüller et al., 2024).
@@
Figure 1: (Left) MaxUCB (blue line) outperforms combined search (black line) to identify the best-performing model class (brown line). (Middle) The irregular distribution of the empirical performance of model classes is left-skewed, and a higher mean may not correspond to a higher maximum. (Right) MaxUCB selects for which model to run one iteration of HPO during two-level optimization.
@@
A popular approach to address the CASH problem is to use categorical and conditional hyperparameters to run HPO directly on the combined hierarchical search space of models and hyperparameters. AutoML systems use this approach, which we call combined search, to search well-performing ML pipelines (Thornton et al., 2013; Feurer et al., 2015; Komer et al., 2014; Kotthoff et al., 2017; Feurer et al., 2022), but HPO remains inefficient in high-dimensional and hierarchical search spaces. A naive solution to address the scalability limitation is to run HPO independently for the smaller search spaces of each ML model class and then compare the found solutions. However, this solution often exceeds available computational resources and does not scale well with an increasing number of ML models. Figure 1 (left) illustrates the difference between searching each space individually (colored dashed lines) and combined search (black line) on an exemplary dataset.
@@
To leverage the efficiency of HPO in low dimensions, we use a Multi-Armed Bandit (MAB) method (Lattimore and Szepesvári, 2020) to dynamically allocate our budget. As shown in Figure 1 (right), each time the bandit strategy pulls an arm, it runs one iteration of HPO to evaluate a new configuration, resulting in a loss. The negative of the loss is used as reward feedback for the bandit algorithm. This approach is known as two-level CASH (decomposed CASH)  (Hoffman et al., 2014; Liu et al., 2019).
@@
While most classical MAB problems aim to maximize the average rewards over time, the goal of the bandit algorithm for two-level CASH should be to maximize the maximum reward observed over time: as illustrated in Figure 1 (middle), tuning the LightGBM (red) will eventually outperform CatBoost. This goal aligns with Max $K$-Armed Bandit (MKB) problems (Carpentier and Valko, 2014; Achab et al., 2017; Baudry et al., 2022), often referred to as Extreme Bandits.
@@
We precisely address this open question through a thorough statistical analysis of the empirical reward distributions of HPO tasks. Our main contribution is a state-of-the-art algorithm for two-level CASH based on a novel extreme bandit algorithm we call MaxUCB (Algorithm 1). We demonstrate the performance of our method on four benchmarks, which highlights the relevance of our assumptions for a wide variety of CASH problems. We analyze the theoretical performance of MaxUCB (Theorem 4.2) through regret bounds that also justify our novel choice of exploration bonus for the type of distributions relevant to the CASH problem. Importantly, our objective is rather to propose a practical algorithm with good empirical performance on CASH rather than a general multi-purpose bandit algorithm, so our guarantees hold under carefully crafted assumptions that resolve previously open questions (Nishihara et al., 2016).
@@ 2 Solving CASH using Bandits
For our approach, we study the decomposed variant (Hoffman et al., 2014; Liu et al., 2019) and address the following two-level optimization problem depicted in Figure 1 (right): at the upper level, we aim to find the overall best-performing ML model $A^{(i^{*})}$ by selecting model $A^{(i)}\in\mathcal{A}$ iteratively, and at the lower level, we aim to find the best-performing configuration $\boldsymbol{\lambda^{*}}\in\boldsymbol{\Lambda}^{(i)}$ for the selected model $A^{(i)}$. Formally,
$A^{(i^{*})}\in\operatorname*{arg\,min}_{A^{(i)}\in\mathcal{A}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}^{*}},{\mathbb{D}}),\quad\text{s.t.}\quad\boldsymbol{\lambda}^{*}\in\operatorname*{arg\,min}_{\boldsymbol{\lambda}\in\boldsymbol{\Lambda^{(i)}}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}},{\mathbb{D}}).$ (2)
The right-hand side of Equation 2 in the lower level can be efficiently addressed by existing iterative HPO methods such as Bayesian optimization (BO) (Jones et al., 1998; Garnett, 2022), which has been demonstrated to perform well in practical settings (Snoek et al., 2012; Chen et al., 2018; Cowen-Rivers et al., 2022). BO fits a surrogate model and uses an acquisition function to find a promising configuration to evaluate next. On the upper level, or left-hand side of Equation 2, the challenge is to carefully allocate the budget $T$ of HPO runs to the $K$ models in a manner that trades off exploration of the hyperparameter space of all models and exploitation (optimization) of the most promising model. As already noted in previous work, this is a typical MAB problem (Cicirello and Smith, 2005; Streeter and Smith, 2006a; Nishihara et al., 2016; Metelli et al., 2022).
At time $t$, the bandit algorithm chooses model $I_{t}\in\mathcal{A}$, and we denote $\boldsymbol{\lambda}_{t}$ the configuration proposed by the HPO method in the lower level. As a reward $r_{i,t}$, we feed back to the bandit algorithm an evaluation of the negative loss:
@@
In general, and as opposed to standard MAB, this reward process is not i.i.d. conditionally on the arm choices because the loss of the models depends on the progress of HPO on each model class (arm) as well as additional loss evaluation noise. To be able to design a tractable bandit algorithm, it is crucial to find an appropriate way to model this process to build controllable estimators. We focus on this aspect in the next section.
To complete the bandit model of this problem, we need to choose a regret metric that defines the oracle objective we compare to, and indeed aligns with Equation 2. For HPO, the regret should target max-value objectives (Jamieson and Talwalkar, 2016; Nishihara et al., 2016).
@@
Instead of using MKB algorithms to directly address Equation 4, related prior works focus on alternative methods. For example, Hu et al. (2021) proposed and analyzed the Extreme-Region Upper Confidence Bounds (ER-UCB) algorithm maximizing the extreme region of the feedback distribution, assuming Gaussian rewards. More recently, Balef et al. (2024) have shown that existing MKB algorithms underperform when applied to two-level CASH, and they proposed methods for maximizing the quantile values instead of the maximum value.
As another alternative method, Li et al. (2020) framed the CASH problem as a Best Arm Identification (BAI) task and introduced the Rising Bandits algorithm (Li et al., 2020). This MAB method assumes that the reward function for each arm increases with each pull, following a rested bandit model with non-decreasing payoffs (Heidari et al., 2016) (which has been shown to have linear regret when the reward increment per pull exceeds a threshold (Metelli et al., 2022)). Rising Bandits can be used for our setting using the maximum observed performance of the HPO history as the reward. However, this algorithm assumes deterministic rewards and increasing concave reward functions. To weaken this assumption, Li et al. (2020) introduced a hyperparameter to increase initial exploration. Mussi et al. (2024) further weakened this assumption by assuming that the moving average of the rewards is an increasing concave function.
Generally, in BAI approaches, the goal is to identify the arm with the highest mean reward. Here, rewards are the result of HPO runs, so this objective does not align well with Equation 2: there is no reason to measure the quality of a model on average over a random subset of hyperparameters chosen by HPO. This approximation made by prior work is justified by the complexity of this modeling problem and the existence of solid foundations on BAI to build upon. But in this work, we propose a fully data-driven model that better fits the true CASH objective and results in better empirical results.
@@ 3 Data Analysis of HPO Tasks
The reward process in Equation 3 is complex as it depends on the model and on the chosen HPO algorithm. So, as discussed, it is necessary to model it. Rather than choosing a convenient parametric family, we conduct a thorough analysis of typical sequences of losses obtained on real benchmarks. For each ML model class (corresponding to arms in our setup), we run $T=200$ iterations of HPO, with $32$ repetitions (each using a different seed) on a varying number of datasets on four AutoML benchmarks (see Appendix D.2 and  C.1).
We first analyze the survival function of the reward distributions. Recall that for a random variable $X\sim d$, the survival function is defined by: $x\mapsto G(x)=P_{X\sim d}(X\geq x).$ Figure 2 shows the average empirical survival function of all observed performances (normalized between 0 and $1$ for each task) for each arm. We rank model classes (i.e., arms) based on their best performance per dataset and report results on two benchmarks (see Appendix C for more details and results on all benchmarks). We observe that the reward distributions are
@@
Figure 2: (Left) The average empirical survival function of the reward distribution per arm ranked per dataset. Thin lines correspond to segments of the reward sequence and show the distribution change over time. (Right) The average empirical survival function per dataset for the best and worst arm. Thin lines correspond to individual datasets.
@@
Suppose Assumption 3.2 holds. Then, there exists $L,U\geq 0$ such that the survival function $G$ can be bounded near $b$ by two linear functions:
$\forall\epsilon\in(0,b-a),\quad L\epsilon\leq G(b-\epsilon)\leq U\epsilon.$ (5)
Lemma 3.3 (proof in Appendix B.1) provides a way to characterize the shape of any bounded distribution near its maximum value through distribution-dependent constants, $L$ and $U$ (see Appendix C.2 for more details and a visualization). Intuitively, it quantifies the behavior of the distribution of the ML model performance in a given hyperparameter space near the optimum. A large value of $L$ indicates a steep drop in the survival function near the maximum, while a small value of $U$ leads to a more gradual decay of the survival function and conversely (see Figure 3).
@@
Figure 4: Histogram of $L$ and $U$ values across individual datasets.
@@
Our empirical analysis shows that existing MKB algorithms’ assumptions about reward distributions are either overly broad or misaligned with the CASH problem. We identify two key differences between our assumptions and those made by current MKB algorithms. First, the sub-optimality gap makes the regret definition (Equation 4) for bounded distributions meaningful. Without the existence of a nonzero gap, the regret definition fails since any policy consistently selecting a single arm can achieve zero regret, as Nishihara et al. (2016) pointed out with Bernoulli distributions. Second, our reward distributions differ from those used in existing MKB algorithms. Lemma 3.3 characterizes the shape of the distribution, with a higher $L$ ensuring more mass near extreme values, making extreme values easier to estimate. In our case, values for $L$ are mostly higher than $1$ while for heavy-tailed distributions, commonly used as the basis for MKB algorithms, are close to 0 (see the rightmost plot in Figure 3). In general, our analysis shows that considering the right range of $L$ values unlocks the problem raised by Nishihara et al. (2016), who focused on constructing counterexamples through the distributions that have unrealistic values for $L$. Thus, under our assumptions, their negative result [^3] Theorem $11$ in (Nishihara et al., 2016): “no policy can be guaranteed to perform asymptotically as well as an oracle that plays the single best arm for a given time horizon.”, which means any policy needs to explore all arms for budget T does not apply here.
@@ 4 MaxUCB
Our exploration bonus $(\frac{\alpha\log(t)}{n})^{2}$ deviates from typical UCB literature due to faster concentration of maximum values in bounded distributions. This is because the probability of bad events (violating confidence intervals for the expectation of the max, see Equation 30 in Appendix B.3) can be written as:
$P(\text{Bad events})\leq O\left(e^{-n\sqrt{C(n)}}+nC(n)\right),$ (6)
where distribution-dependent constants are hidden for clarity. Then, setting $C(n)=1/n^{2}$ minimizes the probability of bad events: the first term becomes independent of $n$, while the second term decreases with $n$. Notably, this faster concentration can only be obtained for the reasonably well-behaved distributions we consider following the study of the previous section and it is not a general property of the maxima; more details can be found in Appendix B.4. Furthermore, MaxUCB uses $\alpha\geq 0$ to control the exploration-exploitation trade-off; a higher $\alpha$ leads to more exploration.
@@
[…]
@@
3:  Pull arm $i$, set $n_{i}\leftarrow 1$, observe reward $r_{i,1}$ $\triangleright$ Evaluating default configuration
@@
7:    Update policy $U_{i}=\max{(r_{i,1},...,r_{i,n_{i}})}+(\frac{\alpha\log(t)}{n_{i}})^{2}$ $\triangleright$ differs from classical UCB, where $U_{i}=\bar{r}_{i}+\sqrt{\tfrac{\alpha\log(t)}{n_{i}}}$
@@
Analysis of MaxUCB. We first show a regret decomposition result specific to max $K$-armed bandits that directly relates the regret definition in equation 4 with the number of suboptimal trials.
@@
(Regret Upper Bound) the regret upper bound up to time $T$ is related to the number of times sub-optimal arms are pulled:
@@
where $N_{i}(T)=\mathbbm{E}(\sum\nolimits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, and $b_{i}$ is the upper bound on the support of the rewards of arm $i$.
@@
$N_{i}(T)\leq\frac{T^{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}}{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}+2\alpha\sqrt{U_{i}T}\log(T).$ (8)
The result of Theorem 4.2 (proof in Appendix B.3) highlights that MaxUCB primarily leverages two key properties: the sub-optimality gap $\Delta_{i}$ and the shape of the distribution, as defined in Lemma 3.3. Specifically, the performance improves with a larger sub-optimality gap $\Delta_{i}$ and higher values of $L_{i^{*}}$ ($L$ for the optimal arm), which means that samples drawn from the distribution of the optimal arm are likely to be close to the extreme values. Additionally, smaller values of $U_{i}$ ($U$ for a sub-optimal arm $i$), which means it is less likely to draw samples close to the extreme values, reduce the number selecting sub-optimal arm $i$, thus enhancing overall performance.
For our task, as is shown in Figure 3, the values for $L_{i^{*}}$ are higher than $1$ and $U_{i}$ less than $10$ in most cases, yielding high empirical performance. We compare the number of pulls observed in our experiments with the expected values based on our theoretical analysis in Appendix E.2, showing that our analysis is not loose. However, it is essential to note that, in general, finding the optimal arm is challenging if $\Delta_{i}$ is close to zero or $L_{i^{*}}$ is very small. We assess the performance of MaxUCB on synthetic tasks in Appendix E.4, showing that the performance of MaxUCB deteriorates on tasks that do not satisfy our assumptions.
Finally, we provide a regret upper bound that combines the decomposition in Proposition 4.1 with the individual upper bounds above. This requires finding a parameter $\alpha$ that resolves a trade-off between the arms and minimizes the total upper bound, as shown in Corollary 9, whose proof is immediate.
@@
If $L_{i^{*}}$, $\min\limits_{i\neq i^{*}}(\Delta_{i})$, and $T$ are known in advance, then the total regret $R(T)$ can be bounded as follows by choosing the exploration parameter $\alpha$ appropriately:
$R(T)\leq\mathcal{O}\bigg(\frac{K\log T}{\sqrt{T}}\max\limits_{i\leq K}b_{i}\bigg),\quad\text{when }\alpha=\frac{1}{4L_{i^{*}}\sqrt{\min\limits_{i\neq i^{*}}\Delta_{i}}}\bigg(1-\frac{2\log(\log(T))}{\log(T)}\bigg).$ (9)